Fermi National Laboratory


Accelerator Update

Wednesday October 8 through Friday October 10
All shifts began in shutdown.

On Wednesday, Operations received a report of a Main Injector quadrupole magnet (Q206) with a LCW leak. It seemed to be an internal leak. Experts will continue to investigate.

On Thursday, technicians began the magnet bake out at Tevatron sector F0. Booster techs completely valved in the Booster LCW system. Operators on the evening shift monitoring the LCW system noticed what might have indicated a leak. They toured the ring, discovered a leak, and valved off the entire system. The leak was due to a hose that had popped off.

ANTIPROTON SOURCE
I need to clear up some inaccuracies I made about the Antiproton Source's ten new motorized quadrupoles stands. The experts had the stands installed in the Debuncher ring (and not in AP2) to provide the best beam positions in the Debuncher. AP2 is upstream of the Debuncher. They hope to increase the amount of antiprotons sent from AP2 into the Debuncher by having better orbit control in the Debuncher. The new stands will control both horizontal and vertical motion. This means that the quadrupole itself can be used as a steering element if the beam doesn't pass through its center. Since tunnels and components can move for various reasons, and orbits are dynamic rather than static, it is likely that these quads will be moved, though infrequently. However, once positioned the experts will turn off control to the motors. The plan is to motorize twenty more Debuncher quadrupoles during next summer's shutdown.

TEVATRON

  • E2: Sector E2 has passed vacuum certification and highpot testing. All that remains for that sector is to unroll 19 magnets and to finish the TeV net suvey.
  • A2: The conning tower work in A2 will be done by today. Next week the crew will move to B3.
  • A48: Techs have installed the new cryo system bypass at sector A48. The bypass has been vacuum certified and its interface connection completed. The collimator and Pots should go in today. As you can see by the pictures below, interfacing is no easy job. It involves soldering and insulating within a three-inch area.
  • Lambertsons: The Lambertson bake-out is underway. The experts expect it to take two days to heat up, five days at temperature, and then two more days to cool off.
  • B3 has an ice-ball in its beam pipe. Techs will warm the house cryo system today and then begin vacuum pumping. Conning tower work will start next week.
  • HLS: The HLS system has been installed and the water should go in today or Monday. Controls people are working on a way to read this out via ACNET. There will be an installation of a cheaper system in B sector to compare performances.
  • The TeV Net survey is going well. The laser tracker crews are done with A, B, C, D, and F sectors and a small part of E sector is underway.
  • Alignment of A2, A4, and E2 should start next week
  • Roll measurements have been completed. Crews are checking data and redoing a few sections. All calibration of the protractors and tilt meters were finished over the weekend with lists generated for magnets rolled more than 1 mr in E2, A2, and A4.
  • Crews have begun cable pulls for the new octupole circuits. The BPM testing is approximately 80% done.
  • Core drilling for 6 new sump pits was completed.
  • The Collider Detector Facility (CDF) wall is done in the Tevatron. The steel inside of CDF will be done in 2 weeks. Techs will monitor the low beta quads with tilt meters during the steel stacking. Alignment crews will conduct a final survey before startup.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 10/10/2003   email Fermilab